OpenStudy (anonymous):

Find the derivative of g(x)= -3/ squareroot x

OpenStudy (anonymous):

\[g(x)=\frac{-3}{\sqrt{x}}=\frac{-3}{x^{\frac{1}{2}}}=-3x^{\frac{-1}{2}}\]Now you can just use the power rule.

OpenStudy (anonymous):

\[\frac{d}{dx}x^{n}=n\times x^{n-1}\]

OpenStudy (anonymous):

for petes sake to not use the product rule with constants

OpenStudy (anonymous):

quotient rule then ?

OpenStudy (anonymous):

no power rule

OpenStudy (anonymous):

@Aylin wrote the method above

OpenStudy (anonymous):

Power rule, Aylin have posted it up here.

OpenStudy (anonymous):

\[y'= -\frac{ 3 }{ 2 }x^{-\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

like that @satellite73 @Aylin

OpenStudy (anonymous):

You didn't execute the power rule properly.

OpenStudy (anonymous):

\[g(x) = -3x ^{\frac{ -1 }{ 2 }}\]g'(x) will be: \[g'(x)=(-3*-\frac{ 1 }{ 2 })*x ^{-\frac{ 1 }{ 2 }-1}=\frac{ 3 }{ 2 } *x^{\frac{ -3 }{ 2 }} \iff g'(x)=\frac{ 3 }{ 2\sqrt{x^3} }\]

OpenStudy (anonymous):

Okay but the thing is you use the power rule first and then take the derivative ?:S

OpenStudy (anonymous):

No, the power rule is a way to differentiate. :]

OpenStudy (anonymous):

@zairhenrique then why is it done twice here ?

OpenStudy (anonymous):

It's only done once

OpenStudy (anonymous):

I am really confused

OpenStudy (anonymous):

the "power rule" is not the same as "write in exponential form" they are two different things

OpenStudy (anonymous):

for example, if you want the derivative of \(f(x)=\frac{2}{x^3}\) then you can REWRITE it as \[f(x)=2x^{-3}\] and then use the power rule to get \[f'(x)=-3\times 2x^{-3-1}=-6x^{-4}\]

OpenStudy (anonymous):

\[g(x) = \frac{ -3 }{ \sqrt{x}} \]

OpenStudy (anonymous):

that is not "using the power rule twice" that is "write in exponential form" then "use the power rule"

OpenStudy (anonymous):

\[g(x)=-3x^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

writing \[g(x)=-3x^{-\frac{1}{2}}\] is not the power rule, it is getting it in the form so that you can use the power rule

OpenStudy (anonymous):

Yes i know

OpenStudy (anonymous):

btw \[\frac{3}{\sqrt{x}}\neq -3x^{\frac{1}{2}}\]it is \[-3x^{-\frac{1}{2}}\]

OpenStudy (anonymous):

ohh i see the problem make sure to write it in exponential form correctly

OpenStudy (anonymous):

\[g'(x)=\frac{ -3 }{ 2 }x^\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

ops that exponent is supposed ot be negative

OpenStudy (anonymous):

\[= \frac{ -3 }{ 2\sqrt{x} }\]

OpenStudy (anonymous):

It's not right... The -3/2 became positive and the x is x³

OpenStudy (anonymous):